sinx=

sinx=

Grade:11

6 Answers

azhar ali
8 years ago
secA+tanA=p then, cosecA=
Forum Team
8 years ago
Dear Rohan,
 
Please post complete question always so that users can answer your question properly.
 
Thanks
Forum Team
SRUTI
8 years ago
secA^2-tanA^2=1from above
secA+tanA=1/secA-tanA
secA+tanA=P ….….….….…...EQN1THEN
secA-tanA=1/P….….….…...EQN2
SOLVING EQN1&2 WE GET
secA=(P^2+1)/2P ,
tanA=(P^2-1)/2P
SINA/COSA=(P^2-1)/2P
SINA*SECA=(P^2-1)/2P
SINA*(P^2+1)/2P =(P^2-1)/2P
SINA=(P^2-1)/(P^2+1)
COSECA=(P^2+1)/(P^2-1)
raju
8 years ago
√(1-cos^2 x)
OUTPASS
8 years ago
√(1-cos^2 x)
SHANMUKESHWAR
8 years ago
√(1-cos^2 x) OR 1/COSECx